OpenStudy (anonymous):

Solve the set of linear equations by Guassian elimination method : a+2b+3c=5, 3a-b+2c=8, 4a-6b-4c=-2. Find c

OpenStudy (anonymous):

Is this algebra 2?

OpenStudy (anonymous):

\[ \begin{bmatrix} 1&2&5\\ 3&-1&2\\ 4&-6&-4 \end{bmatrix} \begin{bmatrix} a\\b\\c \end{bmatrix} = \begin{bmatrix} 5 \\ 8 \\-2 \end{bmatrix} \]